a 1200 kg car travelling at 20 m/s collides with a stationary 1400 kg car. The two cars lock together. Determine the speed of the vehicles immediately after the collision if 53.9% of the initial kinetic energy is converted to heat.
1
Expert's answer
2020-12-06T17:19:56-0500

The speed of the vehicles immediately after the collision can be found from the momentum conservation:


"m_1v_1 + m_2v_2 = (m_1+m_2)v"

where "m_1 = 1200kg, m_2 = 1400kg" are masses of the first and second car respectively, and "v_1 = 20m\/s, v_2 = 0m\/s" are the speeds of the first and second car respectively. "v" is the speed of the vehicles immediately after the collision. Thus, obtain:


"v = v_1\dfrac{m_1}{m_1+m_2} = 20\cdot \dfrac{1200}{1200+1400} \approx 9.2m/s"

Answer. 9.2 m/s.